OpenStudy (anonymous):

8 people are boarding a plane 2 have tickets for first class and board before everyone. In how many ways can these people board the aircraft? Please Explain!!!

OpenStudy (anonymous):

all the 8 people? \[2\times 1\times 6\times 5\times 4\times 3\times 2\times 1\]

OpenStudy (anonymous):

yeah but wouldnt it be different if the first two have to board 1st?

OpenStudy (anonymous):

that is why i put the \(2\times 1\) in front. otherwise it would be 8!

OpenStudy (anonymous):

oh ok thanks!